# Convex Geometry of Building Sets

### v3

@@ 2. Semilattices and Building Sets
Given a meet-semilattice $\mathcal{L}$, we denote $\mathbb{B}(\mathcal{L})=\{B\subseteq\mathcal{L}^{+}:B\leavevmode\nobreak\ \text{is a building set of}\leavevmode\nobreak\ \mathcal{L}\}.$ We will now investigate the poset $(\mathbb{B}(\mathcal{L}),\subseteq)$. When $\mathcal{L}$ is clear from context we may simply refer to $(\mathbb{B}(\mathcal{L}),\subseteq)$ as $\mathbb{B}$. The following observation extends a result of Feichtner-Sturmfels [21, Lemma 3.10] from the Boolean lattice to an arbitrary meet-semilattice.
@@
We first utilize Proposition 2.11 for showing that the intersection of two buildings sets is a building set. Let $B,B^{\prime}\in\mathbb{B}$. As $I(\mathcal{L})\subseteq B,B^{\prime}$, we have that $I(\mathcal{L})\subseteq B\cap B^{\prime}$. Let $x,y\in B\cap B^{\prime}$ such that $x\wedge y\neq\hat{0}$. If $x\vee y\in\mathcal{L}$, then $x\vee y\in B$ and $x\vee y\in B^{\prime}$ as $B$ and $B^{\prime}$ are building sets. Therefore $x\vee y\in B\cap B^{\prime}$. The poset $\mathbb{B}$ is a meet-semilattice with a maximum element $\mathcal{L}^{+}$, hence it is a lattice. ∎
@@ 3. Convex Geometries
Every closure system $(\mathcal{E},\mathbb{S},\sigma)$ comes with a natural meet and join: for $A,B\in\mathbb{S}$, $A\wedge B=A\cap B$ and $A\vee B=\sigma(A\cup B)$. Thus $\mathbb{S}$ is an intersection lattice. Conversely, given an intersection lattice $\mathbb{S}$ with ground set $\mathcal{E}$, we can define a closure operator $\sigma:2^{\mathcal{E}}\rightarrow 2^{\mathcal{E}}$ as follows: given $X\subseteq\mathcal{E}$, let
@@
It was observed by Feicthner-Sturmfels [21, Lemma 3.10] that in the case of the Boolean lattice, Proposition 2.12 implies that the collection of building sets determined a closure operator. Thus far, our article has been concerned with extending results of de Concini-Procesi [14], Postnikov [32], and Feicthner-Sturmfels [21] to Feichtner-Kozlov’s setting of general meet-semilattices [20]. We now go beyond the works of those authors, even in the case of the Boolean lattice, in understanding the structure of building sets.
@@
[16, 25] Let $(\mathcal{E},\mathbb{S},\sigma)$ be a closure system and assume in addition that $\sigma$ satisfies
$\bullet\,\,\text{for every}\leavevmode\nobreak\ A\in\mathbb{S}\leavevmode\nobreak\ \text{and any distinct}\leavevmode\nobreak\ x,y\not\in A,\text{if}\,\leavevmode\nobreak\ y\in\sigma(A\cup\{x\})\,\text{then}\,x\not\in\sigma(A\cup\{y\}).$
@@
Greedoids are a generalization of matroids which were originally introduced by Korte and Lovász in the setting of greedy algorithms [26]. Antimatroids form another important class of greedoids. They first appeared in the work of Dilworth [15] where they were described in the language of lattices. Greeoids, antimatroids, and convex geometries, much like matroids, admit several cryptomorphic definitions. We refer the reader to the excellent survey by Björner and Ziegler for an introduction to these topics [9]. The following definition is one of the most well-known definitions of an antimatroid.
@@
    for $\text{for every}\leavevmode\nobreak\ A,B\in\mathbb{S}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ B\nsubseteq A,\leavevmode\nobreak\ \text{there exists}\leavevmode\nobreak\ e\in B\backslash A\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ A\cup\{e\}\in\mathbb{S}$.
For any set system $(\mathcal{E},\mathbb{S})$, there is a complementary set system $(\mathcal{E},\mathbb{S}^{c})$ where $\mathbb{S}^{c}=\{\mathcal{E}\backslash F:F\in\mathbb{S}\}$. A fundamental result of Björner states that a set system $(\mathcal{E},\mathbb{S})$ is a convex geometry if and only if the complementary set system $(\mathcal{E},\mathbb{S}^{c})$ is an antimatroid [8]. Thus a convex geometry can alternately be defined as follows.
@@
    for $\text{for every}\leavevmode\nobreak\ A,B\in\mathbb{S}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ B\nsupseteq A,\leavevmode\nobreak\ \text{there exists}\leavevmode\nobreak\ e\in A\backslash B\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ A\backslash\{e\}\in\mathbb{S}$
@@
Let $(\mathcal{E},\mathbb{S})$ be a convex geometry then $\mathbb{S}$ is an intersection lattice.
@@
Let $A,B\in\mathbb{S}$. By repeated application of Definition 3.7 condition (2), we may remove all of the elements in $A\setminus B$ from $A$ to produce $A\cap B\in\mathbb{S}$. Therefore, $\mathbb{S}$ is an intersection meet-semilattice with a maximum element $\mathcal{E}$, hence $\mathbb{S}$ is an intersection lattice. ∎
Supersolvable lattices were first introduced by Stanley [34, Definition 1.1], motivated by the study of supersolvable groups. In the same work, he defined supersolvable join-distributive lattices. Edelman proved that join-distributive lattices are cryptomorphic to antimatroids [16, Theorem 3.3]. [^8] More accurately, Edelman proved that meet-distributive lattices are cryptomorphic to convex geometries. Hawrylycz and Reiner demonstrated that the collection of all closure operators on a poset defines a supersolvable join-distributive lattice [24, Theorem 10]. Armstrong provided a natural cryptomorphic description of supersolvable join-distributive lattices in the context of antimatroids, and applied this notion to the study of the sorting order of a Coxeter groups [6].
@@
[6, Definition 2.11] Let $(\mathcal{E},\mathbb{S})$ be a set system. Let $<$ be a total order of $\mathcal{E}$. We say that $\mathbb{S}$ is a supersolvable antimatroid if
@@
    for $\text{for every}\leavevmode\nobreak\ A,B\in\mathbb{S}\leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ B\nsubseteq A,\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ e=\text{min}_{<}(B\backslash A)\leavevmode\nobreak\ \text{we have }\leavevmode\nobreak\ A\cup\{e\}\in\mathbb{S}.$
We say $\mathbb{S}$ is supersolvable with respect to $<$.
Thus one has the following dual definition of supersolvable convex geometries, which was utilized by Gillespie in the study of matroid activities [23, Proposition 15].
@@
[6] [23] [^9] Definition 3.10 and the phrase “supersolvable convex geometry” do not appear in [6, 23], nonetheless we feel it is appropriate to attribute this definition to those sources as it is implicit in their discussions. Let $(\mathcal{E},\mathbb{S})$ be a set system. Let $<$ be a total order of $\mathcal{E}$. We say that $\mathbb{S}$ is a supersolvable convex geometry if
@@
    for $\text{for every}\leavevmode\nobreak\ A,B\in\mathbb{S}\leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ B\nsupseteq A,\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ e=\text{min}_{<}(A\backslash B)\leavevmode\nobreak\ \text{we have }\leavevmode\nobreak\ A\setminus\{e\}\in\mathbb{S}.$
We say $\mathbb{S}$ is supersolvable with respect to $<$.
De Concini-Procesi proved the following statement in the case of subspace arrangements, which was generalized by van der Veer to meet-semilattices [^10] Van der Veer utilized this result for proving that the motivic zeta function of a hyperplane arrangement is independent of the building set..
@@
[14, Section 2.5][35, Lemma 2] Let $\mathcal{L}$ be a meet-semilattice and $B\in\mathbb{B}(\mathcal{L})$. Let $<$ be a linear extension of $\mathcal{L}$, then for $x=\text{min}_{<}(B\backslash I(\mathcal{L}))$ we have $B\setminus\{x\}\in\mathbb{B}(\mathcal{L})$.
We offer the following extension of Proposition 3.11. We note that by Proposition 3.8, the closed sets of a convex geometry form an intersection lattice, hence Proposition 3.12 subsumes Proposition 2.12.
@@
Let $<$ be a linear extension of $\mathcal{L}$. As was noted in example 2.8, $\mathcal{L}^{+}\in\mathbb{B}$ thus condition (1) is satisfied. For verifying condition (2), let $B,B^{\prime}\in\mathbb{B}$ with $B^{\prime}\not\subseteq B$, and $x=\text{min}_{<}(B^{\prime}\backslash B)$. We claim that $B^{\prime}\backslash\{x\}\in\mathbb{B}$, and we verify the conditions of Proposition 2.11. Since $I(\mathcal{L})\subseteq B$ and $x\not\in B$, $x$ is not irreducible so $B^{\prime}\backslash\{x\}$ contains $I(\mathcal{L})$. Let $y,z\in B^{\prime}\backslash\{x\}$ such that $y\wedge z\neq\hat{0}$. If $y\vee z\in\mathcal{L}$, by Proposition 2.11, $y\vee z\in B^{\prime}$. We wish to show that $y\vee z\neq x$ for which we need only consider $y,z<x$. By the minimality of $x$, we know $y,z\in B$. By Proposition 2.11, $y\vee z\in B$ so certainly $y\vee z\neq x$. Thus $y\vee z\in B^{\prime}\backslash\{x\}$. By Proposition 2.11, it follows that $B^{\prime}\backslash\{x\}\in\mathbb{B}$. ∎
@@
This a standard property of any convex geometry – we provide a proof here. For $B\in\mathbb{B}(\mathcal{L})$, set $B_{\circ}:=B\backslash I(\mathcal{L})$. Define the function $r:\mathbb{B}\rightarrow\mathbb{Z}_{\geq 0}$ by $r(B)=|B_{\circ}|$, and note that $r(I(\mathcal{L}))=0$. Let $B,B^{\prime}\in\mathbb{B}$ such that $B^{\prime}$ covers $B$. By Proposition 3.12, if $x=\text{min}_{<}(B^{\prime}\setminus B)$, then $B^{\prime}\backslash\{x\}\in\mathbb{B}$, thus $B^{\prime}=B\cup\{x\}$, and $r(B^{\prime})=r(B)+1$. Therefore $\mathbb{B}(\mathcal{L})$ is ranked. To see that $r$ is supermodular, let $B,B^{\prime}\in\mathbb{B}$. Then $r(B\vee B^{\prime})=|(B\vee B^{\prime})_{\circ}|\geq|B_{\circ}\cup B^{\prime}_{\circ}|$ and $r(B\wedge B^{\prime})=|(B\wedge B^{\prime})_{\circ}|=|(B\cap B^{\prime})_{\circ}|=|B_{\circ}\cap B^{\prime}_{\circ}|$ together imply
@@
Let $\sigma:2^{\mathcal{E}}\rightarrow 2^{\mathcal{E}}$ be a closure operator and $<$ be a total order on $\mathcal{E}$. We say that $\sigma$ is a supersolvable closure operator (with respect to $<$) if for each closed set $A$ and $e\in\mathcal{E}\setminus A$, we have that
@@
Suppose that $(\mathcal{E},\mathbb{S})$ is a supersolvable convex geometry with associated closure operator $\sigma$. Suppose that $\sigma$ is not supersolvable, then there exists a closed set $A$ and $e,f\in\mathcal{E}\setminus A$ such that $f\in\sigma(A\cup\{e\})\setminus(A\cup\{e\})$, and $f<e$. Take $f$ to be the minimum such element with respect to $<$. Applying condition (2) of Definition 3.10 to the pair of sets $A,\sigma(A\cup\{e\})$, we have that $\sigma(A\cup\{e\})\setminus f\in\mathbb{S}$, but this contradicts the definition of $\sigma(A\cup\{e\})$.
@@
Let $\mathcal{L}$ be a meet-semilattice and $X\subset\mathcal{L}^{+}$. Set $X_{0}:=X\cup I(\mathcal{L})$. Given $X_{i}$, set $Y_{i}:=\{a\vee b\in\mathcal{L}:a,b\in X_{i}\,,\,a\wedge b\neq\hat{0}\}$ and $X_{i+1}:=X_{i}\cup Y_{i}.$ Let $k$ be such that which $X_{k+1}=X_{k}$, then $\sigma(X)=X_{k}$.
@@
Because $\mathcal{L}$ is finite and $X_{i}\subseteq X_{i+1}$, there must be some $k$ such that $X_{k+1}=X_{k}$. By Proposition 2.11, it follows that $X_{k}$ is a building set, so $\sigma(X)\subseteq X_{k}$. We establish the opposite inclusion by induction. Clearly, $X_{0}\subseteq\sigma(X)$. Suppose from some $i$ we have that $X_{i}\subseteq\sigma(X)$, then $Y_{i}\subseteq\sigma(X)$, hence $X_{i+1}\subseteq\sigma(X)$. ∎
We describe a method of producing building sets. First take some $X,X^{\prime}\subseteq\mathcal{L}^{+}$. Construct $\sigma(X),\sigma(X^{\prime})$ using Proposition 3.18. If $\sigma(X)\nsubseteq\sigma(X^{\prime})$ We can then apply supersolvability to remove elements from $\sigma(X)$ one at a time – all of the intermediate sets will be building sets.
@@ 4. Restrictions of building sets and nested set complexes
In this section we utilize the perspective developed in this article for proving Theorem 4.5 which states that buildings sets and nested set complexes behave well with respect to certain restrictions of meet-semilattices. Our theorem is inspired by the following wonderful result of Mantovani-Padrol-Pilaud [28] for oriented matroids, which is motivated by fundamental polyhedral considerations. [^12] We remark that the recent work of Brauner-Eur-Pratt-Vlad [12] has some overlap with the work of Mantovani-Padrol-Pilaud [28], but not in this particular aspect it seems.
@@
[28] Suppose that $(E,\mathcal{M})$ is a simple acyclic oriented matroid with Las Vergnas’ face lattice $\mathcal{F}$, and let $\mathcal{B}_{n}=(2^{E},\subseteq)$ be the Boolean lattice where $|E|=n$. Let $\mathcal{F}\hookrightarrow\mathcal{B}_{n}$ be the natural embeddining. Then
@@
Our primary goal in proving Theorem 4.5 is to establish an unoriented matroid version of Theorem 4.1, which has a separate polyhedral motivation coming from Hodge theory for matroids.131313 We originally learned the statement of Theorem 4.1 from Padrol’s lecture at the Simons Center for Geometry and Physics Workshop on “Combinatorics and Geometry of Convex Polyhedra” in March 2023. In personal communication, we asked Mantovani-Padrol-Pilaud whether the natural unoriented version of Theorem 4.1 should hold, and they said that they had considered, but not investigated this question. After an earlier draft of this work appeared on arXiv, Matt Larson contacted us to ask whether the unoriented version of Theorem 4.1 (now Corollary 4.6) follows from the results of this paper. Inspired by this independent interest, the authors verified this statement which resulted in Theorem 4.5 – we thank Matt Larson for his question. We thank Arnau Padrol and Vincent Pilaud for their encouragement with the result of this section, and for giving us their blessing to present it here. Although we will not define an oriented matroid, which is not our focus here, we will now define a matroid. Because the perspective of this article is lattice theoretic, we provide the definition of a matroid in terms of its lattice of flats.
@@
Let $M$ be a matroid on ground set $E$ with $|E|=n$. Let $\mathcal{L}$ be the lattice of flats of $M$ and let $B$ be a building set on $\mathcal{L}$. The Bergman fan of $M$ with respect to $B$ is the following collection of polyhedral cones in $\mathbb{R}^{n}\,$:
@@
Wonderful compactifications of the torus (the complement of the coordinate arrangement) are smooth projective toric varieties, and nestohedral fans are the complete fans associated to these varieties. [^14] This correspondence holds after first quotienting the Bergman fan by its lineality space $\langle{\bf 1}\rangle_{\mathbb{R}}$. Direct combinatorial proofs that nestohedral fans are smooth and projective are given in Postnikov [32, Proposition 7.10] and Zelevinsky [36, Corollary 5.2]. Feichtner-Sturmfels [21, Proposition 3.6] prove that for arbitrary $M$ and $B$, the Bergman fan $\Sigma$ forms a unimodular, but typically incomplete, fan whose support is independent of the choice of $B$. The Bergman fan with respect to the maximum building set, as introduced in Ardila-Klivans [5], is a subfan of the braid arrangement, i.e. the nestohedral fan with respect to the maximum building set. Several researchers have presented interesting nested set complex structures for Bergman fans as restrictions of nestohedral fans: Braden-Huh-Matherne-Proudfoot-Wang describe the augmented Bergman fan as a subfan of the stellahedral fan [11, Proposition 2.6], Crowley-Huh-Larson-Simpson-Wang describe the polymatroid Bergman fan as a subfan of the polypermutahedral fan [13, Definition 1.7], and Eur-Larson describe the augmented Bergman fan of a polymatroid as a subfan of the polystellahedral fan [19, Definition 3.7].

### v4

@@ 2. Semilattices and Building Sets
We will now investigate the poset $(\mathbb{B}(\mathcal{L}),\subseteq)$. When $\mathcal{L}$ is clear from context we may simply refer to $(\mathbb{B}(\mathcal{L}),\subseteq)$ as $\mathbb{B}$. The following observation extends a result of Feichtner-Sturmfels [feichtner2005matroid, Lemma 3.10] from the Boolean lattice to an arbitrary meet-semilattice.
@@
We first utilize Proposition 2.11 for showing that the intersection of two buildings sets is a building set. Let $B$ and $B^{\prime}$ be building sets. As $I(\mathcal{L})\subseteq B,B^{\prime}$, we have that $I(\mathcal{L})\subseteq B\cap B^{\prime}$. Let $x$ and $y$ be elements of $B\cap B^{\prime}$ such that $x\wedge y\neq\hat{0}$. If $x\vee y\in\mathcal{L}$, then $x\vee y\in B$ and $x\vee y\in B^{\prime}$ as $B$ and $B^{\prime}$ are building sets. Therefore $x\vee y\in B\cap B^{\prime}$. The poset $\mathbb{B}$ is a meet-semilattice with a maximum element $\mathcal{L}^{+}$, hence it is a lattice. ∎
@@ 3. Convex Geometries
[…]
@@
It was observed by Feichtner-Sturmfels [feichtner2005matroid, Lemma 3.10] that in the case of the Boolean lattice, Proposition 2.12 implies that the collection of building sets determined a closure operator. Thus far, our article has been concerned with extending results of de Concini-Procesi [de1995wonderful], Postnikov [postnikov2009permutohedra], and Feichtner-Sturmfels [feichtner2005matroid] to Feichtner-Kozlov’s setting of general meet-semilattices [feichtner2004incidence]. We now go beyond the works of those authors, even in the case of the Boolean lattice, in understanding the structure of building sets.
@@
[edelman1980meet, jamison1982perspective] Let $(\mathcal{E},\mathbb{S},\sigma)$ be a closure system and assume in addition that $\sigma$ satisfies
$\bullet\,\,\text{for every}\penalty 10000\ A\in\mathbb{S}\penalty 10000\ \text{and any distinct}\penalty 10000\ x,y\not\in A,\text{if}\,\penalty 10000\ y\in\sigma(A\cup\{x\})\,\text{then}\,x\not\in\sigma(A\cup\{y\}).$
@@
Greedoids are a generalization of matroids which were originally introduced by Korte and Lovász in the setting of greedy algorithms [korte1981mathematical]. Antimatroids form another important class of greedoids. They first appeared in the work of Dilworth [dilworth1940lattices] where they were described in the language of lattices. Greedoids, antimatroids, and convex geometries, much like matroids, admit several cryptomorphic definitions. We refer the reader to the excellent survey by Björner and Ziegler for an introduction to these topics [bjorner1992introduction]. The following definition is one of the most well-known definitions of an antimatroid.
@@
    $\text{for every}\penalty 10000\ A,B\in\mathbb{S}\penalty 10000\ \text{such that}\penalty 10000\ B\nsubseteq A,\penalty 10000\ \text{there exists}\penalty 10000\ e\in B\backslash A\penalty 10000\ \text{such that}\penalty 10000\ A\cup\{e\}\in\mathbb{S}$.
[…]
@@
    $\text{for every}\penalty 10000\ A,B\in\mathbb{S}\penalty 10000\ \text{such that}\penalty 10000\ B\nsupseteq A,\penalty 10000\ \text{there exists}\penalty 10000\ e\in A\backslash B\penalty 10000\ \text{such that}\penalty 10000\ A\backslash\{e\}\in\mathbb{S}$
@@
If $(\mathcal{E},\mathbb{S})$ is a convex geometry then $\mathbb{S}$ is an intersection lattice.
@@
Let $A,B\in\mathbb{S}$. By repeated application of Definition 3.7 condition (2), we may remove all of the elements in $A\setminus B$ from AA to produce $A\cap B\in\mathbb{S}$. Therefore, $\mathbb{S}$ is an intersection meet-semilattice with a maximum element $\mathcal{E}$, hence $\mathbb{S}$ is an intersection lattice. ∎
Supersolvable lattices were first introduced by Stanley [stanley1972supersolvable, Definition 1.1], motivated by the study of supersolvable groups. In the same work, he defined supersolvable join-distributive lattices. Edelman proved that join-distributive lattices are cryptomorphic to antimatroids [edelman1980meet, Theorem 3.3]. [^8] More accurately, Edelman proved that meet-distributive lattices are cryptomorphic to convex geometries. Hawrylycz and Reiner demonstrated that the collection of all closure operators on a poset defines a supersolvable join-distributive lattice [hawrylycz1993lattice, Theorem 10]. Armstrong provided a natural cryptomorphic description of supersolvable join-distributive lattices in the context of antimatroids, and applied this notion to the study of the sorting order of a Coxeter group [armstrong2009sorting].
@@
[armstrong2009sorting, Definition 2.11] Let $(\mathcal{E},\mathbb{S})$ be a set system. Let << be a total order of $\mathcal{E}$. We say that $\mathbb{S}$ is a supersolvable antimatroid if
@@
    $\text{for every}\penalty 10000\ A,B\in\mathbb{S}\penalty 10000\ \text{with}\penalty 10000\ B\nsubseteq A,\penalty 10000\ \text{for}\penalty 10000\ e=\text{min}_{<}(B\backslash A)\penalty 10000\ \text{we have }\penalty 10000\ A\cup\{e\}\in\mathbb{S}.$
We say $\mathbb{S}$ is supersolvable with respect to <<.
Thus one has the following dual definition of supersolvable convex geometries, which was utilized by Gillespie in the study of matroid activities [gillespie2020convexity, Proposition 15].
@@
[armstrong2009sorting] [gillespie2020convexity] [^9] Definition 3.10 and the phrase “supersolvable convex geometry” do not appear in [armstrong2009sorting, gillespie2020convexity], nonetheless we feel it is appropriate to attribute this definition to those sources as it is implicit in their discussions. Let $(\mathcal{E},\mathbb{S})$ be a set system. Let << be a total order of $\mathcal{E}$. We say that $\mathbb{S}$ is a supersolvable convex geometry if
@@
    $\text{for every}\penalty 10000\ A,B\in\mathbb{S}\penalty 10000\ \text{with}\penalty 10000\ B\nsupseteq A,\penalty 10000\ \text{for}\penalty 10000\ e=\text{min}_{<}(A\backslash B)\penalty 10000\ \text{we have }\penalty 10000\ A\setminus\{e\}\in\mathbb{S}.$
We say $\mathbb{S}$ is supersolvable with respect to <<.
De Concini-Procesi proved the following statement in the case of subspace arrangements, and van der Veer showed the result for meet-semilattices [^10] Van der Veer utilized this result for proving that the motivic zeta function of a hyperplane arrangement is independent of the building set..
@@
[…]
Feichtner-Müller established the following strengthened version.
Proposition 3.12.
[feichtner2005topology, prooof of Theorem 4.2] Let $\mathcal{L}$ an atomic meet-semilattice, and $B,B^{\prime}\in\mathbb{B}(\mathcal{L})$ with $B^{\prime}\subsetneq B$. Let << be a linear extension of $\mathcal{L}$, then for $x=\text{min}_{<}(B\backslash B^{\prime})$ we have $B\setminus\{x\}\in\mathbb{B}(\mathcal{L})$.
We offer the following extension of Proposition 3.12. We note that by Proposition 3.8, the closed sets of a convex geometry form an intersection lattice, hence Proposition 3.13 subsumes Proposition 2.12. This result can be equivalently interpreted as an extension Proposition 3.12 to the setting where $B$ and $B^{\prime}$ may not be comparable, or as a combination of Proposition 3.12 with Proposition 2.12.
@@
[…]
@@
[…]
@@
Let $\sigma:2^{\mathcal{E}}\rightarrow 2^{\mathcal{E}}$ be a closure operator and << be a total order on $\mathcal{E}$. We say that $\sigma$ is a supersolvable closure operator (with respect to <<) if for each closed set $A$ and $e\in\mathcal{E}\setminus A$, we have that
@@
Suppose that $(\mathcal{E},\mathbb{S})$ is a supersolvable convex geometry with associated closure operator $\sigma$. Suppose that $\sigma$ is not supersolvable, then there exists a closed set $A$ and $e,f\in\mathcal{E}\setminus A$ such that $f\in\sigma(A\cup\{e\})\setminus(A\cup\{e\})$, and $f<e$. Take $f$ to be the minimum such element with respect to <<. Applying condition (2) of Definition 3.10 to the pair of sets $A,\sigma(A\cup\{e\})$, we have that $\sigma(A\cup\{e\})\setminus f\in\mathbb{S}$, but this contradicts the definition of $\sigma(A\cup\{e\})$.
@@
Let $\mathcal{L}$ be a meet-semilattice and $X\subset\mathcal{L}^{+}$. Set $X_{0}:=X\cup I(\mathcal{L})$. Given $X_{i}$, set $Y_{i}:=\{a\vee b\in\mathcal{L}:a,b\in X_{i}\,,\,a\wedge b\neq\hat{0}\}$ and $X_{i+1}:=X_{i}\cup Y_{i}.$ Let $k$ be such that $X_{k+1}=X_{k}$, then $\sigma(X)=X_{k}$.
Note that because $\mathcal{L}$ is finite and $X_{i}\subseteq X_{i+1}$, there exists some $k$ such that $X_{k+1}=X_{k}$.
@@
By Proposition 2.11, it follows that $X_{k}$ is a building set, so $\sigma(X)\subseteq X_{k}$. We establish the opposite inclusion by induction. Clearly, $X_{0}\subseteq\sigma(X)$. Suppose from some $i$ we have that $X_{i}\subseteq\sigma(X)$, then $Y_{i}\subseteq\sigma(X)$, hence $X_{i+1}\subseteq\sigma(X)$. ∎
We briefly describe a general method to produce building sets. First, take some $X,X^{\prime}\subseteq\mathcal{L}^{+}$. Construct $\sigma(X),\sigma(X^{\prime})$ using Proposition 3.19. If $\sigma(X)\nsubseteq\sigma(X^{\prime})$, we can apply supersolvability to remove elements from $\sigma(X^{\prime})\setminus\sigma(X)$ one at a time. All of the intermediate objects will be building sets.
@@ 4. Restrictions of building sets and nested set complexes
In this section we utilize the perspective developed in this article for proving Theorem 4.5 which states that buildings sets and nested set complexes behave well with respect to certain restrictions of meet-semilattices. Our theorem is inspired by the following wonderful result of Mantovani-Padrol-Pilaud [mantovani2025facial] for oriented matroids, which is motivated by fundamental polyhedral considerations. [^12] We remark that the recent work of Brauner-Eur-Pratt-Vlad [brauner2024wondertopes] has some overlap with the work of Mantovani-Padrol-Pilaud [mantovani2025facial], but not in this particular aspect it seems.
@@
[mantovani2025facial] Suppose that $(E,\mathcal{M})$ is a simple acyclic oriented matroid with Las Vergnas’ face lattice $\mathcal{F}$, and let $\mathcal{B}_{n}=(2^{E},\subseteq)$ be the Boolean lattice where $|E|=n$. Let $\mathcal{F}\hookrightarrow\mathcal{B}_{n}$ be the natural embeddining. Then
@@
Our primary goal in proving Theorem 4.5 is to establish an unoriented matroid version of Theorem 4.1, which has a separate polyhedral motivation coming from Hodge theory for matroids. [^13] We originally learned the statement of Theorem 4.1 from Padrol’s lecture at the Simons Center for Geometry and Physics Workshop on “Combinatorics and Geometry of Convex Polyhedra” in March 2023. In personal communication, we asked Mantovani-Padrol-Pilaud whether the natural unoriented version of Theorem 4.1 should hold, and they said that they had considered, but not investigated this question. After an earlier draft of this work appeared on arXiv, Matt Larson contacted us to ask whether the unoriented version of Theorem 4.1 (now Corollary 4.6) follows from the results of this paper. Inspired by this independent interest, the authors verified this statement which resulted in Theorem 4.5 – we thank Matt Larson for his question. We thank Arnau Padrol and Vincent Pilaud for their encouragement with the result of this section, and for giving us their blessing to present it here. Update from arXiv version 4: The work of Mantovani-Padrol-Pilaud is now available [mantovani2025facial]. Although we will not define an oriented matroid, which is not our focus here, we will now define a matroid. Because the perspective of this article is lattice theoretic, we provide the definition of a matroid in terms of its lattice of flats.
@@
Let $M$ be a matroid on ground set $E$ with $|E|=n$. Let $\mathcal{L}$ be the lattice of flats of $M$ and let $B$ be a building set on $\mathcal{L}$ with $B\in E$ [^14] Some authors do not require building sets for matroids to contain the maximum element when discussing Bergman fans (see [eur2025building, mantovani2025facial]). Our result readily extends to that level of generality.. The Bergman fan of $M$ with respect to $B$ is the following collection of polyhedral cones in $\mathbb{R}^{n}\,$:
@@
Wonderful compactifications of the torus (the complement of the coordinate arrangement) are smooth projective toric varieties, and nestohedral fans are the complete fans associated to these varieties. [^15] This correspondence holds after first quotienting the Bergman fan by its lineality space $\langle{\bf 1}\rangle_{\mathbb{R}}$. Direct combinatorial proofs that nestohedral fans are smooth and projective are given in Postnikov [postnikov2009permutohedra, Proposition 7.10] and Zelevinsky [zelevinsky2006nested, Corollary 5.2]. Feichtner-Sturmfels [feichtner2005matroid, Proposition 3.6] prove that for arbitrary $M$ and $B$, the Bergman fan $\Sigma$ forms a unimodular, but typically incomplete, fan whose support is independent of the choice of $B$. The Bergman fan with respect to the maximum building set, as introduced in Ardila-Klivans [ardila2006bergman], is a subfan of the braid arrangement, i.e. the nestohedral fan with respect to the maximum building set. Several researchers have presented interesting nested set complex structures for Bergman fans as restrictions of nestohedral fans: Braden-Huh-Matherne-Proudfoot-Wang describe the augmented Bergman fan as a subfan of the stellahedral fan [braden2022semi, Proposition 2.6], Crowley-Huh-Larson-Simpson-Wang describe the polymatroid Bergman fan as a subfan of the polypermutahedral fan [crowley2022bergman, Definition 1.7], and Eur-Larson describe the augmented Bergman fan of a polymatroid as a subfan of the polystellahedral fan [eur2023intersection, Definition 3.7].